OpenStudy (anonymous):

how would i find the least common denominator 7/12 - 1/9 ?

OpenStudy (anonymous):

factorize denominators, 12=2*2*3 9=3*3 common denominator must have all terms, also 2*2*3*3

OpenStudy (osanseviero):

12 * 1 = 12 12 * 2 = 24 12 * 3 = 36 9 * 1 = 9 9 * 2 = 18 9 * 3 = 27 9 * 4 = 36 Both in 36 For step by step use http://www.helpwithfractions.com/least-common-denominator.html

OpenStudy (anonymous):

36

OpenStudy (anonymous):

@mik_wind and osanseviero -> thank you, this definatly helps

OpenStudy (anonymous):

@reheen->thank you

OpenStudy (anonymous):

yw